OpenStudy (anonymous):

Solve the differential equation (2x+7y)dx-(4x-y)dy=0 .... Thanks ! =)

OpenStudy (lgbasallote):

use homogeneous thingies

OpenStudy (anonymous):

can you show me how ? and how did you know it's homogenous ?

OpenStudy (anonymous):

homogeneous thingies !

OpenStudy (anonymous):

it is non Exact differential equation. you will have to find integrating factor to make it exact.

OpenStudy (anonymous):

yes. thank you. but can you explain more ? how did you know it's homogenous and how can i solve it ? thanks !

OpenStudy (anonymous):

you can use x=tx and y=ty if you can take out t back then it is homogeneous.

OpenStudy (anonymous):

@lgbasallote help him .sorry i got a go :(

OpenStudy (lgbasallote):

normally when everything has the same degree it is homogeneous...

OpenStudy (lgbasallote):

andeverything here has degree of 1 so i know it is homogeneous

OpenStudy (lgbasallote):

bye everything i mean all terms

OpenStudy (unklerhaukus):

\[(2x+7y)\text dx-(4x-y)\text dy=0\]\[2x+7y\text dx=(4x-y)\text dy\]\[\frac{2x+7y}{4x-y}=\frac{\text dy}{\text dx}\] \[\frac{\text dy}{\text dx}=\frac{2x+7y}{4x-y}=\frac{2+7\frac yx}{4-\frac yx}\] let \[\qquad v=\frac yx\qquad y=vx\qquad \frac {\text dy}{\text dx}=v+x\frac{\text dv}{\text dx}\] \[v+x\frac{\text dv}{\text dx}=\frac{2+7v}{4-v}\]

OpenStudy (unklerhaukus):

which is variables separable ,

OpenStudy (anonymous):

thank you very very much !!!!